### v1

# PTEB: Towards Robust Text Embedding Evaluation via Stochastic Paraphrasing at Evaluation Time with LLMs

Abstract
Current evaluations of sentence embedding models typically rely on static test beds such as the Massive Text Embedding Benchmark (MTEB). While invaluable, repeated tuning on a fixed suite can inflate reported performance and obscure real-world robustness. We introduce the Paraphrasing Text Embedding Benchmark (PTEB), a dynamic protocol that stochastically generates meaning-preserving paraphrases at evaluation time and aggregates results across multiple runs. Using a cost-efficient LLM-based method grounded in semantic textual similarity gold ratings, we show that LLMs generate token-diverse but semantically preserving, paraphrases. Across 7 MTEB tasks, we validate our hypothesis that the performance of sentence encoders is sensitive to changes in token space even when semantics remain fixed. We also observe that smaller models are not disproportionately affected relative to larger ones. Our results are statistically robust over multiple runs and we extended our experiments to 3 multilingual datasets covering 10 languages. More generally, we aim to propose a new evaluation paradigm in NLP that relies less on static, pre-defined benchmarks but shifts towards dynamic, stochastic evaluation leveraging eval-time compute.
PTEB: Towards Robust Text Embedding Evaluation via Stochastic Paraphrasing at Evaluation Time with LLMs
Manuel Frank Department of Computer Science Munster Technological University Manuel.Frank@zohomail.eu          Haithem Afli Department of Computer Science Munster Technological University Haithem.Afli@mtu.ie
1 Introduction
Text embeddings have become a cornerstone of Natural Language Processing (NLP), enabling the representation of semantic meaning in dense vector spaces that can be applied to a wide range of downstream tasks. Traditionally, embeddings are evaluated on static benchmarks like the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023), which has become the de facto standard.
Figure 1: PTEB paraphrases data at evaluation time to measure embedding model performance.
More recently, MTEB has been extended to the Massive Multilingual Text Embedding Benchmark (MMTEB) (Enevoldsen et al., 2025), covering more than 250 languages, and RTEB [^1] https://huggingface.co/blog/rteb, a retrieval benchmark with private test sets.
However, benchmarks like MTEB need to be maintained and further developed (Chung et al., 2025). A persistent challenge is that static benchmarks saturate over time. Although this saturation partly reflects genuine progress, it can also be attributed to overfitting. This arises from at least two factors (Liang et al., 2025): 1. Data contamination which refers to the inclusion of benchmark items within training data, enabling models to exploit prior exposure rather than demonstrate authentic modelling capabilities (also known as data leakage). 2. Biased overtraining which involves deliberate allocation of training resources to the domains of anticipated benchmarks, resulting in models with uneven performance that excel in benchmarked domains while underperforming elsewhere. Overall, more robust embedding benchmarks are needed (Goel et al., 2025). So far, the only solutions have been temporary through private test sets or the release of new or updated benchmarks.
To address this issue, we argue that model capabilities should be evaluated across dynamically generated variations of evaluation tasks at evaluation time rather than predefined static benchmarks. We therefore introduce the Paraphrasing Text Embedding Benchmark (PTEB), which generates data variations at evaluation time, creating a "population" of semantically equivalent but textually distinct problem instances that better approximates the diversity of real-world applications. Whereas MTEB targets maximal multilingual breadth and task coverage, PTEB is a complementary, compute-bounded statistically robust protocol that stress-tests semantic invariance via multi-item stochastic eval-time paraphrasing.
Generative LLMs have already been shown to be effective for data augmentation through paraphrasing. In Dai et al. (2023), for example, paraphrases were used to train a BERT model (Devlin et al., 2019). Earlier, Wahle et al. (2022) demonstrated that generative LLMs can paraphrase text in a way that is difficult for humans to identify. Frank and Afli (2024) and Thirukovalluru and Dhingra (2025) used generative models to generate variations of the textual input, e.g. by paraphrasing, which are used for data augmentation at runtime to improve the performance of embeddings.
Prior research on robustness and benchmarking has mainly relied on algorithmic transformations, masked language modelling (MLM), or classical word embedding techniques. For example, nlpaug (Ma, 2019) provides a general-purpose augmentation toolkit with methods such as back-translation, synonym replacement, and embedding-based substitutions (e.g., using Word2Vec (Mikolov et al., 2013)), while TextAttack (Morris et al., 2020) systematises adversarial attacks through synonym swaps, character edits, and MLM substitutions, supporting both benchmarking and adversarial training. Building on these ideas, TextFlint (Wang et al., 2021b) introduced a multilingual robustness framework combining universal and task-specific transformations incl. adversarial attacks, all validated through human evaluation. More recently, Yang et al. (2023) published an article on benchmark contamination and evasion of contamination detection methods. They demonstrated the fragility of existing evaluations for generative models, showing that rephrased benchmark samples can bypass standard decontamination methods and artificially inflate the performance of a 13b parameter model to GPT-4 levels, thus undermining benchmark trustworthiness.
Whereas Yang et al. (2023) use paraphrasing to diagnose weaknesses of static benchmarks for generative LLMs, we employ eval-time paraphrasing to construct inherently contamination-resistant evaluations for embedding models. Earlier work has therefore emphasised robustness testing, augmentation, or contamination detection, but remains fundamentally constrained by static transformation rules, MLM, or classical embeddings. In contrast, our contribution is to harness state-of-the-art generative LLMs as general paraphrasers at eval-time, yielding a stochastic, dynamically moving benchmark that continuously produces novel paraphrastic test variants. This extends robustness testing from predefined perturbations to generatively rich, stochastic evaluation runs, specifically targeting embedding models across diverse NLP tasks. More broadly, our goal is to shift the focus of the NLP community from static datasets and offline augmentation to stochastic evaluation-time benchmarks powered by modern generative models.
The PTEB architecture is visualised in Figure 1.
Dataset Word count
BIOSSES (Soğancıoğlu et al., 2017) 22.8
SICK-R (Marelli et al., 2014) 9.6
STS12 (Agirre et al., 2012) 10.9
STS13 (Agirre et al., 2013) 8.9
STS14 (Agirre et al., 2014) 9.2
STS15 (Agirre et al., 2015) 10.6
STS17 (Cer et al., 2017) 8.7
STS22 (Chen et al., 2022) 461.0
STSB (Cer et al., 2017) 9.9
Average 61.1
Table 1: Average word count per input text for English STS datasets.
2 Method
Figure 2: First, MTEB STS data is used to evaluate different generative LLMs for their STS rating capability. Then, using the selected LLM judge, generative LLMs are evaluated for paraphrasing MTEB STS data. Finally, the best paraphrase model is used to paraphrase MTEB datasets at eval-time generating a dynamic benchmark for embedding models.
Our method is illustrated in Figure 2. We employ generative LLMs in two distinct roles: First, as an LLM judge to select a paraphrase model. Second, as a paraphraser generating paraphrases during evaluation. To evaluate and select the LLM judge, we use sentence pairs from the MTEB STS datasets. In the absence of human evaluations or direct validation of the generated paraphrases, our method anchors PTEB in scientific rigor by grounding it in STS gold scores. Table 1 lists the STS datasets that we use and shows that the STS datasets cover a range of different text lengths ranging from short sentences to STS22’s paragraphs with an average length of 461 words. Throughout this paper, cosine similarity is employed as the similarity metric between embeddings. Generally, we use the same metrics as MTEB, for example, Spearman’s rank correlation for STS.
1. LLM Judge Evaluation. The MTEB STS datasets provide similarity ratings on bounded scales, obtained through annotation methods like Amazon Mechanical Turk. We evaluate generative LLMs based on their ability to assess the semantic similarity of sentence pairs vs. these STS gold ratings in order to select an LLM judge model.
2. Paraphrase Model Evaluation. Next, we apply generative LLMs to paraphrase the MTEB STS datasets. The LLM judge selected in the previous step rates the semantic similarity of the paraphrases on a scale from 0 to 5. Based on semantic similarity, edit distance (ED), and runtime, we select the paraphrase model ("paraphraser"). We include the ED to avoid textually overly similar paraphrases compared to the original text (larger ED is better).
3. Embedding Model Evaluation on PTEB. Finally, we employ the selected paraphraser to rephrase MTEB datasets at eval-time generating a dynamic benchmark. This allows us to evaluate various embedding models on the paraphrases and compare their scores on the original MTEB datasets. (For non-STS tasks, we use the datasets and metrics listed in Appendix A.)
3 Experiments
In this section, we present the results of our experiments. We begin by outlining the implementation details followed by the results of the LLM judge evaluation. We then examine how the selected LLM judge rates various generative models for paraphrasing. Finally, we present the results of running PTEB across different embedding models. To analyse how well the results generalise to different tasks, we evaluate datasets across all MTEB tasks and include non-English languages.
Figure 3: LLM judge mean average error (MAE) and number of examples per gold score on MTEB STS.
3.1 Implementation
For the LLM judge and paraphraser selection, we assessed the generative open-weights models gemma 3-27b (Gemma Team et al., 2025), gpt-oss-20b (OpenAI et al., 2025), and qwen3-32b (Yang et al., 2025). The implementation details incl. hyperparameters can be found in Appendix B.
As encoders, we evaluated qwen3-8b (Zhang et al., 2025), mxbai-embed-large-v1 (Lee et al., 2024), e5-mistral-7b-instruct (Wang et al., 2024a, b), and embeddinggemma-300m (Vera et al., 2025) as SOTA encoders. As a baseline, we included the Sentence Transformer all-mpnet-base-v2 (Reimers and Gurevych, 2020). These models were selected to include both models that currently lead or rank high on the MTEB leaderboard and a baseline model with lower performance. To ensure consistency between original and paraphrased conditions, the evaluation pipeline was re-implemented following the MTEB specification in Muennighoff et al. (2023). [^2] Our scripts are not identical with the current MTEB-evaluation but follow the original paper, e.g. MTEB now undersamples classification data to include 8 examples per label instead of using the complete train/test splits. For hyperparameters see Appendix B; for all prompts Appendix C.
3.2 Experimental Results
3.2.1 LLM Judge Evaluation
Analysing the mean absolute error (MAE) relative to the gold rating distribution (that is, the similarity scores) shows that all models have lower errors on sentence pairs with high ($\geq 4$) similarity scores Figure 3). However, while the Qwen3 and the GPT-OSS models have lower MAEs on sentence pairs with label close to 0, this does not hold for the Gemma-3 model. Table 2 presents the results for each LLM judge across datasets. Overall, gemma-3-27b achieved the highest performance on MTEB STS. With the exception of STS22 the model outperformed GPT-OSS and Qwen3 on all datasets.
Examining performance by dataset across word count bins, Figure 4 shows that sentences with $\geq 20$ words yield the lowest performance (average score 64.49%), while sentences with <10<10 words achieve an average Spearman’s rank correlation of $\geq 80\%$. In particular, the models show lower performance on long sentence pairs in STS14 and STSB with scores of 31.59% and 54.26%, respectively.
Figure 4: LLM judge scores on MTEB STS for different word count intervals; empty cells indicate that the dataset does not contain sentences with the respective word count. (in %)
Dataset gemma-3-27b gpt-oss-20b qwen-3-32b
BIOSSES 88.53 83.75 80.94
SICK-R 76.61 70.84 72.21
STS12 75.13 66.95 70.42
STS13 89.12 87.07 87.60
STS14 83.82 81.95 81.55
STS15 89.87 87.26 86.70
STS17 91.42 88.99 87.53
STS22 66.87 68.82 70.59
STSB 87.26 84.50 85.67
Average 83.18 80.02 80.89
Table 2: LLM judge scores on MTEB STS; best scores bold. (in %)
STS Judge Model Size
270m 1b 4b 12b 27b
1.9 33.28 72.74 82.76 83.18
Table 3: LLM judge ablation on the numbers of parameters for Gemma-3 grading paraphrase semantic similarity on the STS datasets. (in %)
To assess the effect of model size, we performed an ablation by varying the number of parameters (Table 3). Gemma-3 12b performs comparably to the 27b model (82.76% vs. 83.18%). By contrast, the 4b and, especially, the 1b and 270m variants exhibit substantial performance drops (72.74%, 33.28%, and 1.9% respectively).
3.2.2 Paraphrase Model Evaluation
Using gemma-3-27b as the LLM judge, we observe that also as a paraphraser it not only achieves semantic similarity scores close to the best-performing paraphraser, gpt-oss-20b, but produces paraphrases with larger ED at a lower runtime (Table 4). Hence, overall, gemma-3-27b is the most suitable paraphraser for PTEB.
3.2.3 Embedding Model Evaluation on PTEB
As single‐run estimates can be unreliable (Reimers and Gurevych, 2018; Dror et al., 2019), we evaluate STS with $n=6$ different paraphrasing seeds. Across datasets, all models decline on PTEB vs. originals with the smallest drop for embeddinggemma-300m (Table 5). e5-mistral-7b-instruct decreases by 2.70%, while others (except Gemma) drop by up to 5.09%. The gap is smallest on STS22 and largest on STS12.
Also the standard deviation across runs differs substantially between datasets. On SICK-R and STS15 the sample standard deviation ($\sigma$) is at the lower end ($\sigma\leq 0.09$ and $\sigma\leq 0.20$ respectively) in contrast to BIOSSES ($\sigma\in[1.01,2.23]$).
Paraphrase
Model
Semantic
Sim.
Edit
Distance
Run-
time
gemma-3-27b 4.23 0.81 0.27
gpt-oss-20b 4.41 0.55 1.24
qwen3-32b 4.36 0.63 5.35
Table 4: Paraphrase evaluation results showing average semantic similarity (0–5 scale) and normalized edit distance compared to originals (in %), as well as runtime (seconds per text) on STS datasets; best results bold.
Figure 5: STSB paraphrase examples.
Dataset
all-mpnet-
base-v2
embedding
gemma-300m
mxbai-embed-
large-v1
e5-mistral-
7b-instruct
qwen3-em-
bedding-8b
Orig. PTEB Orig. PTEB Orig. PTEB Orig. PTEB Orig. PTEB
BIOSSES 80.39 77.22 70.27 64.67 86.09 84.06 84.61 77.68 86.29 81.44
$\Delta$-3.17/±2.23 $\Delta$-5.60/±1.61 $\Delta$-2.03/±1.11 $\Delta$-6.93/±1.01 $\Delta$-4.85/±1.53
SICK-R 80.60 77.45 71.99 70.28 82.79 78.24 80.76 79.01 84.97 81.88
$\Delta$-3.15/±0.09 $\Delta$-1.71/±0.07 $\Delta$-4.55/±0.07 $\Delta$-1.75/±0.05 $\Delta$-3.09/±0.07
STS12 72.64 61.85 63.53 54.95 79.07 67.91 75.83 68.23 81.41 67.21
$\Delta$-10.79/±0.29 $\Delta$-8.58/±0.41 $\Delta$-11.16/±0.24 $\Delta$-7.60/±0.32 $\Delta$-14.20/±0.30
STS13 83.49 78.91 67.71 73.70 89.81 84.19 84.31 84.35 87.89 82.90
$\Delta$-4.58/±0.56 $\Delta$+5.99/±0.64 $\Delta$-5.62/±0.33 $\Delta$+0.04/±0.35 $\Delta$-4.99/±0.41
STS14 78.00 72.14 64.95 65.03 85.22 78.41 79.26 78.12 83.32 77.10
$\Delta$-5.86/±0.26 $\Delta$+0.08/±0.19 $\Delta$-6.81/±0.12 $\Delta$-1.14/±0.21 $\Delta$-6.22/±0.22
STS15 85.66 81.58 77.19 74.47 89.34 85.05 85.20 86.20 89.56 86.50
$\Delta$-4.08/±0.11 $\Delta$-2.72/±0.19 $\Delta$-4.29/±0.20 $\Delta$+1.00/±0.10 $\Delta$-3.06/±0.12
STS17 90.61 85.06 82.61 75.28 89.22 85.00 88.86 85.61 92.19 88.60
$\Delta$-5.55/±0.25 $\Delta$-7.33/±0.70 $\Delta$-4.22/±0.28 $\Delta$-3.25/±0.31 $\Delta$-3.59/±0.21
STS22 68.31 69.39 58.05 63.21 69.02 68.97 69.66 67.49 69.98 68.61
$\Delta$+1.08/±0.48 $\Delta$+5.16/±0.67 $\Delta$-0.05/±0.38 $\Delta$-2.17/±0.74 $\Delta$-1.37/±0.66
STSB 83.43 74.90 67.45 66.97 89.29 82.20 84.58 82.02 88.50 84.56
$\Delta$-8.53/±0.35 $\Delta$-0.48/±0.53 $\Delta$-7.09/±0.36 $\Delta$-2.56/±0.38 $\Delta$-3.94/±0.32
Average 80.35 75.39 69.31 67.62 84.43 79.34 81.45 78.75 84.90 79.87
$\Delta$-4.96 $\Delta$-1.69 $\Delta$-5.09 $\Delta$-2.70 $\Delta$-5.03
Table 5: Embedding model performance on original and PTEB STS datasets (in %). Each cell shows the original and PTEB score; the difference is denoted by $\Delta$. Red indicates performance decreases; green improvements. All values are aggregated over $n=6$ runs with the sample standard deviation over the different paraphrases denoted by ±. Bold marks row-wise best for original and paraphrased text separately.
To test statistical significance, we use the paired Wilcoxon signed-rank test, which can be applied in many NLP settings (Dror et al., 2018). When pooling over all STS datasets, all embedding models except for embeddinggemma-300m outperform the baseline all-mpnet-base-v2 significantly after Holm-correction (Holm, 1979) ($p\ll\alpha=0.05$; for all models: $0\notin\text{CI}_{95\%}(\Delta_{HL})$; see Table 6). [^3] The Holm-method adjusts p-values when testing multiple hypotheses. It sorts the $m$ hypotheses by their p-values ($p_{i}$) and then checks, starting with the lowest p-value, for a given significance level $\alpha$ if $\forall i\in\{1,\dots,m\}:p_{i}<\frac{\alpha}{m-i+1}$ (Calonico and Galiani, 2025). The estimated location shifts $\Delta_{HL}$ [^4] The Hodges–Lehmann estimator (Hodges and Lehmann, 1963) associated with the Wilcoxon signed-rank test is the median of all Walsh averages, i.e., the pairwise means formed from every self-pair and unique unordered pair of the observed within-pair differences. are negative for all models except embeddinggemma-300m.
mpnet vs. $\Delta_{HL}$ [CI] p-value
embedding- +7.33 6.50e-10
gemma-300m [7.01, 8.12]
mxbai-embed- -3.74 1.33e-08
large-v1 [-5.11, -3.09]
e5-mistral- -3.48 1.40e-07
7b-instruct [-4.37, -2.60]
qwen3-em- -4.53 1.48e-09
bedding-8b [-4.82, -4.19]
Table 6: Comparisons on PTEB STS vs. all-mpnet-base-v2. $\Delta_{HL}$ is the Hodges–Lehmann estimator (Hodges and Lehmann, 1963) with its 95% confidence interval; negative values favour the listed model. P-values are Holm-adjusted. ($n=6$ runs)
3.2.4 Ablation on STS Embedding Model Size
To analyse the influence of the encoder model size, we conduct an ablation study using the Qwen3 models (Table 7). The performance gap between the original and paraphrased datasets remains similar across all three Qwen3 variants, with the smallest model even displaying the narrowest gap.
STS Score Embedding Model Size
0.6b 4b 8b
Original 81.42 84.26 84.90
PTEB 77.19 78.72 79.87
Runtime [s] 135 201 263
VRAM [GB] 1.1 7.6 14.1
Table 7: Qwen3 embedding ablation on the numbers of parameters for STS; STS scores in %. Runtime is the time for encoding only. VRAM refers to memory usage of the model only. Best in bold. ($n=6$ runs)
Task
all-mpnet-
base-v2
embedding
gemma-300m
mxbai-embed-
large-v1
e5-mistral-
7b-instruct
qwen3-em-
bedding-8b
Orig. PTEB Orig. PTEB Orig. PTEB Orig. PTEB Orig. PTEB
Classification 92.23 85.21 85.83 80.56 93.66 87.42 90.99 84.56 92.78 86.64
Banking77 $\Delta$ -7.02 $\Delta$ -5.27 $\Delta$ -6.24 $\Delta$ -6.43 $\Delta$ -6.14
Clustering 50.18 48.24 32.99 34.33 51.25 48.93 45.61 39.04 55.51 49.46
TwentyNewsGroups $\Delta$ -1.94 $\Delta$ +1.34 $\Delta$ -2.32 $\Delta$ -6.57 $\Delta$ -6.05
Pair Classif. 73.85 71.36 59.03 63.73 78.54 73.11 75.36 73.78 72.22 73.90
TwitterSemEval $\Delta$ -2.49 $\Delta$ +4.70 $\Delta$ -5.43 $\Delta$ -1.58 $\Delta$ +1.68
Reranking 65.82 64.42 48.14 48.20 65.20 63.98 59.96 59.51 67.48 65.27
AskUbuntuDupQ. $\Delta$ -1.40 $\Delta$ +0.06 $\Delta$ -1.22 $\Delta$ -0.45 $\Delta$ -2.21
Retrieval 45.89 45.45 29.66 32.89 65.11 61.18 53.52 49.89 75.27 73.62
ArguAna $\Delta$ -0.44 $\Delta$ +3.23 $\Delta$ -3.93 $\Delta$ -3.63 $\Delta$ -1.65
Summarisation 26.25 23.74 20.35 15.74 35.55 32.59 34.24 32.80 38.87 35.29
SummEval $\Delta$ -2.51 $\Delta$ -4.61 $\Delta$ -2.96 $\Delta$ -1.44 $\Delta$ -3.58
Average 59.04 56.40 46.00 45.91 64.88 61.20 59.95 56.60 67.02 64.03
$\Delta$ -2.64 $\Delta$ -0.09 $\Delta$ -3.68 $\Delta$ -3.35 $\Delta$ -2.99
Table 8: Embedding model performance for non-STS tasks and datasets on original data and PTEB. Each cell shows the original and PTEB score; the difference is denoted by $\Delta$. Performance decreases red; improvements green. Bold: best per column. See Appendix A for metrics per dataset and references. ($n=1$ runs; in %)
Model Classification Pair Class. STS
Original PTEB Original PTEB Original PTEB
paraphrase-multilingual-mpnet-base-v2 88.54 85.55 89.14 89.29 83.46 79.98
embeddinggemma-300m 90.51 86.11 86.95 86.49 61.89 59.74
multilingual-e5-large-instruct 89.08 86.32 87.62 87.23 72.82 70.86
qwen3-embedding-8b 89.83 87.04 87.35 87.24 83.35 80.24
Avg. word count text 1/text 2 19.9 18.8 30.5/8.2 30.3/9.4 8.3/8.1 8.8/8.4
Table 9: Multilingual eval: AmazonCounterFactuals (class.), RTE3 (pair class.), STS17 (STS) ($n=6$ runs; in %). Performance drops red; improvements green. Bold: best per column. For paired data, we report both word counts.
Paraphrase Prompt
PTEB STS
Score
Edit
Distance
Default 74.92 80.87
Variant 1 74.24 81.49
Variant 2 74.82 82.00
Variant 3 75.16 83.63
Table 10: PTEB STS scores and the corresponding edit distances to the original text for the default paraphrase prompt and 3 differently phrased prompt variants; see Appendix C.3 for the prompts used. (Based on all-mpnet-base-v2 embeddings; $n=1$ runs; in %)
3.2.5 PTEB for Non-STS Tasks
Extending PTEB to the remaining 6 MTEB tasks, shows that the average performance of all models drops on PTEB compared to original texts (Table 8). qwen3-8b is the best-performing model on 5 out of the 6 tasks. The total performance drop of embeddinggemma-300m is marginal (-0.09) while the other models drop between 2.64 and 3.35. (For an additional visualisation see Appendix Figure 6.)
3.2.6 Multilingual PTEB
We validate our results on non-English data, using RTE3 (Giampiccolo et al., 2007), AmazonCounterFactuals (O’Neill et al., 2021), and the multilingual version of STS17; covering 10 languages (see Table 9 for results; see footnote 6 for a list of language pairs). We used the same models as before if they are natively multilingual, or their multilingual variants alternatively. We find significant performance drops on AmazonCounterFactuals and STS17 while the results are stable on RTE3.
3.2.7 STS Paraphrase Prompting Sensitivity
Since LLMs can be sensitive to prompt changes (Mizrahi et al., 2024), we tested the influence of 3 variations of the paraphrase prompt and compared the PTEB STS scores and the ED between original and paraphrased texts (Table 10). We find limited differences ($\leq 0.68\%$ for the score and $\leq 2.76\%$ for the ED).
4 Discussion
Our method demonstrates that the 3 generative LLMs evaluated in our study are well-suited to rate semantic similarity with STS scores above 80% (Table 2). Gemma-3-27b even outperformed 3 of the 5 embedding models on average on the original STS data (see Table 2 vs. Table 5). [^5] With the limitation that Gemma-3-27b has been evaluated only on one run and is hence not a like-for-like comparison. Furthermore, the evaluated LLMs generate paraphrases with high semantic similarity while introducing token-level variation from the original text (Table 4). We found that gemma-3-27b generated more textually diverse paraphrases at higher throughput: average ED 0.81 vs. 0.55 and 0.63 at a 5× to 20× speed-up compared to gpt-oss-20b and qwen3-32b respectively.
Moreover, across the 7 PTEB tasks and 10 languages, most of the 18 datasets show decreased performance of embedding models on paraphrased data (Table 5, Table 8, Table 9). Our results support the hypothesis of some benchmark saturation and overfitting. While qwen3-8b achieves the highest overall scores on the original data and PTEB, reordering occurs at the dataset level. qwen3-8b, for example, performs best on 4 PTEB datasets where it does not achieve the highest score on the original data, and vice versa; it ranks highest on 3 original datasets where it does not lead on PTEB. This demonstrates that embedding models show different degrees of robustness towards paraphrasing.
For STS22, the relatively small performance drops on PTEB may be explained by its higher word count (avg. 461 per text vs. 61 for all STS data; see Table 1). Based on qualitative observations, we suggest that STS22’s longer texts may allow paraphrases with fewer lexical changes, while paraphrases of shorter sentences rely more on synonym replacement. Moreover, on shorter sentences single tokens have more weight and may distort the embedding model. Another outlier is RTE3 with only marginal score changes. This might stem from its intuitively more robust entailment relation between a long text 1 and a short text 2 (Table 9).
Furthermore, English datasets with initial scores ($\geq 80$), such as BIOSSES, SICK-R, STS13/15/17/B, and Banking77 showed relatively large performance gaps (many models -3 to -6%; also see Appendix Figure 6 which visualises this on task level). This suggests that the initially high scores may depend on shortcuts in token space which are neutralised by paraphrasing. It shows that PTEB’s new evaluation method adds a stress test for semantic invariance, mitigating concerns that MTEB may favour lexical shortcuts and overfitted embedding models. However, neither do we claim that any specific model is more overfitted than another one nor do we make general claims on the performance of individual models as the aim of our study is to present a novel evaluation method; not a comprehensive benchmarking study.
Unexpectedly, smaller encoders were not more sensitive to paraphrasing: Our ablation study on encoder size showed the similar performance and robustness of qwen-3-4b vs. the 8b model with reduced runtime and VRAM (Table 7). For the datasets in our study, the results indicate that robust performance under paraphrase stress can be achieved with smaller embedding models. Furthermore, the prompt sensitivity analysis showed that PTEB is robust to prompt changes further mitigating potential concerns that the results might be brittle due to the use of generative LLMs (Table 10).
5 Conclusion
We introduce PTEB, a sentence embedding evaluation method that stochastically paraphrases MTEB data at eval-time. We show that the paraphrase LLMs yield semantically preserving yet token-diverse outputs. Since our end-to-end LLM-based method indirectly grounds the LLMs in human gold ratings, it can serve as a template for researchers who cannot include expensive human evaluations. Building on this, we demonstrate that the average performance of the tested embedding models drops across 7 MTEB tasks when paraphrased, demonstrating that PTEB effectively uncovers generalisation failures. As a multilingual check, we included 3 multilingual datasets covering 10 languages and 3 different tasks. Ablations over LLM judges and embedding models with different parameter counts show clear runtime–quality trade-offs for the models in scope, and that smaller models (generative LLMs: 12B; sentence encoders: 4B) perform comparably to larger ones on our tests. Our experiments confirm that PTEB yields statistically robust results across multiple runs - even if the paraphrase prompt is varied. In general, PTEB is a novel and more robust benchmarking methodology well suited to augment current embedding benchmarks.
Acknowledgments
PTEB builds upon MTEB (Muennighoff et al., 2023) and MMTEB (Enevoldsen et al., 2025), released under the Apache License 2.0. PTEB is independent work and is not affiliated with or endorsed by the original MTEB or MMTEB authors.
We thank the MTEB maintainers and contributors for the benchmark suite and for making the code publicly available as well as the developers from Hugging Face and of Sentence Transformers for enabling open research.
We also thank the developers of the open-weight models used in this work, including Gemma-3 (Gemma Team et al., 2025; Vera et al., 2025), Qwen3 (Yang et al., 2025), and GPT-OSS (OpenAI et al., 2025).
We acknowledge the use of AI-assisted coding (ChatGPT with GPT-5; Claude Code with Sonnet 4.0/4.5 and Opus 4.1).
6 Limitations
The main limitation of our work lies in the restriction of non-STS datasets to one per task and a single evaluation run. This choice was deliberate: our objective was to demonstrate the applicability of our approach and to provide initial insights rather than to present a comprehensive benchmark of embedding models like MTEB does. Moreover, our analyses of STS datasets using $n=6$ runs already demonstrated the statistical robustness of PTEB.
For the same reason, we limited our evaluation on PTEB to five embedding models and 3 multilingual datasets. In future work, we are planning to implement PTEB as an actual benchmark and include more embedding models and datasets. This is not to replace MTEB but to provide additional signal for embedding model performance.
Furthermore, we did not conduct human evaluations to validate the generated paraphrases. Our goal in this methodological work is an end-to-end LLM-based protocol in which paraphrase models are judged by LLMs calibrated to STS gold ratings, providing an indirect human anchor. This design democratises evaluation by enabling researchers and practitioners with limited budgets to run robust studies without costly, labour-intensive human annotation, while acknowledging that targeted human audits remain valuable for high-stakes settings. Nevertheless, going forward, we plan to include human evaluations to further validate our LLM judge-based approach.
Moreover, the evaluations of the LLM judge and paraphrase models were limited to a single run. This was sufficient for our goal of demonstrating an end-to-end methodology for PTEB that grounds generative model selection in scientific rigor. For production scenarios, our method can serve as a template and be expanded accordingly. Alternatively, one might pragmatically select a paraphrase model directly.
7 Ethical Considerations
We did not identify major ethical concerns arising from this work. However, LLMs (and, hence, PTEB) may replicate or amplify biases present in their training data (Bender et al., 2021). To support ethical scrutiny and reproducibility, we use only open-weight generative models and only embedding models available via Sentence Transformers (i.e., no commercial APIs), enabling researchers to reproduce results, pin versions, and audit behaviour including limited inspection of internals without API accessibility limitations or usage restrictions. While it does not by itself resolve bias, it improves assessability for ethical research.
No personal data was collected and all evaluations are based on public datasets.
Appendix A Non-STS Datasets and Metrics
The following English non-STS datasets (evaluation metrics) were used:
  • Classification: Banking77 (Casanueva et al., 2020) and AmazonCounterFactuals (O’Neill et al., 2021) (Accuracy)
  • Clustering: TwentyNewsGroups Lang (1995) (V-measure)
  • Pair Classification: TwitterSemEval (Xu et al., 2015) and RTE3 Giampiccolo et al. (2007) (average precision)
  • Reranking: AskUbuntuDupQuestions (Wang et al., 2021a) (MAP)
  • Retrieval: ArguAna (Boteva et al., 2016) (nDCG@10)
  • Summarisation: SummEval Fabbri et al. (2021) (Spearman correlation)
Further details on all datasets are available in Muennighoff et al. (2023) and Enevoldsen et al. (2025), as well as on https://huggingface.co/mteb/datasets.
Appendix B Implementation Details
Embedding models. All embedding models were ran using the sentence-transformers package (Reimers and Gurevych, 2020) (v3.3.1), dtype torch.bfloat16 and default max_seq_len (ranging from 384 for all-mpnet-base-v2 to 32,768 for qwen3-embedding-8B). For instruction-tuned models we use the model-specific prompts if available. Please note that some models like qwen3-embedding-8b might achieve better scores with custom prompts. Further information on the models is available at https://huggingface.co/models.
Also, for task-specific hyperparameters we applied the MTEB-defaults (v1.38.33), e.g. max_iter=100 for the Logistic Regression model in the Classification task. For a description of the evaluation protocol for each task the reader may refer to Muennighoff et al. (2023). MTEB is licensed under the Apache License 2.0.
Generative Models. All generative models were run using Ollama (v0.11.7):
  • qwen3:32b
    • Quantisation: Q4_K_M
    • License: Apache License Version 2.0, January 2004
  • gpt-oss:20b
    • Quantisation: MXFP4
    • License: Apache License Version 2.0, January 2004
  • gemma3:27b
    • Quantisation: Q4_K_M
    • License: Gemma Terms of Use (last modified 2024-02-21)
We used the default hyperparameters and environment variables (e.g. OLLAMA_FLASH_ATTENTION=1). Further information on the models is available at www.ollama.com/models.
Random Seeds. We used 1337 as a random seed for all experiments. If more than one run was performed, the seed was incremented by one after each run for the paraphrase models only. An exception are the paraphrases of BIOSSES and SICK-R. For these two no paraphrase have been generated with seed=1342 but seed=1443 instead.
Hardware. All experiments were conducted on a personal computer with the following specification:
  • GIGABYTE GeForce RTX 5090 GAMING OC 32G (GPU)
  • AMD Ryzen 9 9950X (CPU)
  • DDR5-6400 64GB (RAM)
Appendix C Prompts
The following prompts were used for generative LLMs. Curly brackets indicate placeholders that are filled in with the corresponding variables. These prompts were designed in initial experiments using dummy sentences with the goal to provide the desired output. They are not optimised systematically. The generated paraphrases were selectively reviewed by the authors (in multiple, but not all languages).
C.1 LLM Judge Prompts
Used in step 1 of our method to rate STS:
Semantic Similarity Rating Prompt You are an expert evaluator of semantic textual similarity. Your task is to rate how similar two sentences are in meaning. Please rate the semantic similarity between the following two sentences on a scale from 0.0 to 5.0 using one decimal place: – 0.0-1.0: Completely different meanings – 1.1-2.0: Slightly related but mostly different – 2.1-3.0: Somewhat similar with some shared meaning – 3.1-4.0: Very similar with minor differences – 4.0-5.0: Essentially the same meaning Only respond with a single decimal number from 0.0 to 5.0 (one decimal place). Do not provide explanations or additional text. Sentence 1: {sentence1} Sentence 2: {sentence2} Similarity rating:
C.2 Paraphrase Evaluation Prompts
Paraphrase Rating Prompt You are an expert evaluator of semantic textual similarity. Your task is to rate how similar two sentences are in meaning. Please rate the semantic similarity between the following two sentences on a scale from 0.0 to 5.0 using one decimal place: – 0.0-1.0: Completely different meanings – 1.1-2.0: Slightly related but mostly different – 2.1-3.0: Somewhat similar with some shared meaning – 3.1-4.0: Very similar with minor differences – 4.0-5.0: Essentially the same meaning Original: {original} Paraphrase: {paraphrase} Only respond with a single decimal number from 0.0 to 5.0 (one decimal place). Do not provide explanations or additional text. Similarity rating:
C.3 Paraphrase Prompts
Used in step 2 and step 3 of our method to generate paraphrases:
Default Rephrase the following text while keeping its original meaning. Only reply with the paraphrased text and only provide a single response — no alternatives! Do not include thinking tokens, explanations or notes.
Used for multilingual datasets to maintain consistency between input and output language:
Multilingual Rephrase the following text while keeping its original meaning. Only reply with the paraphrased text and only provide a single response — no alternatives! Do not include thinking tokens, explanations or notes. Answer in the same language as the input text. Do not translate to English or any other language.
Used in step 3 of our method to test sensitivity to prompt changes. The variants were obtained by asking GPT-5 to rephrase the default prompt:
Variation 1 Paraphrase the text below so that the meaning stays the same. Return only one rewritten version of the text. Do not add explanations, reasoning steps, or multiple options.
Variation 2 Rewrite the given text in different words but preserve its meaning. Output exactly one paraphrase. Do not include commentary, reasoning, or alternative suggestions.
Variation 3 Provide a single paraphrased version of the following text. The meaning must remain unchanged. Do not include any notes, thoughts, or more than one response.
Appendix D Non-English Languages
The following language pairs are included in the multilingual datasets. Since AmazonCounterFactuals is not a paired dataset, only a single language is given. All languages are encoded in Alpha-2 Country Codes. [^6] https://www.iban.com/country-codes
AmazonCounter-
Factuals
RTE3 STS17
EN DE–DE EN–EN
DE EN–EN AR–AR
JP FR–FR EN–AR
IT–IT EN–DE
EN–TR
ES–EN
ES–ES
FR–EN
IT–EN
KR–KR
NL–EN
Table 11: Languages in AmazonCounterfactuals, RTE3, and STS17.
Appendix E Additional Analyses
Figure 6: Scores per task on original MTEB data vs. PTEB (English only, for original MTEB data based on $n=1$ runs, for PTEB based on $n=6$ runs; in %).